# Monte Carlo simulations of 2D flat-sheet membrane filters for constant-pressure water purification

### v1

@@ Abstract
Membrane filtration is widely used in wastewater treatment to remove foulants from contaminated water. Foulant build-up on the membrane occludes the area open for fluid flow, which impairs the efficiency of the filtration operation by decreasing the flux through the membrane. This necessitates the introduction of cleaning interventions to regenerate the membrane after fouling. Backwashing is a strategy to restore the membrane, wherein clean water is processed backward through the membrane to dislodge attached foulants. We develop a Monte Carlo model based on physical parameters to simulate constant-pressure forward filtration and backwashing through dead-end, flat-sheet membranes, with membrane fouling dominated by intermediate blocking. We validate our model against real-world experiments and show good agreement between experimental results and numerical simulations. The newly developed model can be used to further investigate the impact of varying backwashing duration, frequency, and/or pressure on the rate of flux recovery.
@@ 1 Introduction
Increases in population, industry, and pollution over the past century have exacerbated the long-extant problem of clean water scarcity [12, 77, 79]. Water purification requires many stages of treatment, among which membrane filtration is favorable for its relatively low energy requirements and capital cost [26, 10]. This technology has been applied in many industries, including water treatment for wastewater [8, 74, 59, 59], surface water [75, 36, 54], and groundwater [75, 10]; food and drink manufacturing [70, 15]; and downstream processing in the pharmaceutical industry [56, 76, 14, 62].
Membrane filtration operates primarily through a simple size exclusion mechanism: fluid fouled by suspended solids, bacteria, and other undesired materials is sieved through a semi-permeable barrier whose pores admit the passage of smaller particles and block the passage of larger particles [9]. Membrane filtration processes involve two repeated steps: forward filtration (FF) followed by backwashing (BW). In forward filtration with dead-end, flat-sheet membranes, the fluid is pushed perpendicularly through the membrane and foulants accumulate directly onto its surface. The result is a phenomenon known as membrane fouling, wherein the membrane pores become increasingly blocked [71, 13]. When filtration is run under constant-pressure conditions, membrane fouling results in considerable flux decline, which, in the long run, impairs the efficiency of and decreases the yield from filtration procedures [32, 33]. To counteract this decline, operators employ “backwashing,” running clean water backward through the membrane to dislodge foulants that have attached to it. This process recovers some, though not all, of the flux lost during forward filtration.
@@ 2.2.1 Advection-Diffusion
Our Monte Carlo method builds on equation (3). We define the concentration of a given foulant in the contaminated water being filtered as $F_{\textnormal{conc}}$ (cells  $\text{\,}{\mathrm{m}}^{-3}$) and let $F_{0}$ be the number of foulants in the channel at time $t=0$ based on $F_{conc}$ [49]. We compute $n_{exp}$ (cell  $\text{\,}{\mathrm{s}}^{-1}$) as the number of foulants that enter the domain at every time step based on the initial flux, $J_{0}$ ( $\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$).
@@ 2.2.2 Foulant grafting & flux decline
When a foulant particle approaches the membrane, it either grafts to its surface or it bounces back into the channel flow. Whether each foulant particle attaches or bounces back is determined probabilistically. Any given foulant has an associated probability of grafting to the membrane, $P_{gr}$, that is uniformly distributed between 0 and some maximum probably of grafting $p_{gr}$, i.e., $P_{gr}\in\textnormal{Unif}(0,p_{gr})$. The probability $P_{gr}$ may be additionally adjusted by the particle size and whether other foulants are already stuck on the area of the membrane where a foulant would attach. We allow multiple foulants to occupy the same position on the membrane to conceptualize foulants stacking on top of one another, which in 3D would correspond to having distinct $z$-coordinates.
We model fouling dynamics as dominated by “intermediate pore blocking” where foulants cover both the membrane surface and the membrane pores due to their size [13, 83]; this creates a decline in flux over the course of forward filtration. Experimental analysis suggests that membrane filtration transitions from the intermediate pore blocking regime to the “cake filtration” regime when 10 – 20% of the membrane area remains open [83]. The rate of membrane area coverage depends on the contamination of the water, the membrane pore sizes, and the duration of filtration. At the concentration levels, pore sizes, and filtration runtimes modeled in the experimental set-ups we compare to in Section 3.1, this degree of membrane area coverage is not achieved and, thus, the cake formation phase of filtration is not reached [83].
@@
We adapt equation (4) to model the flux decline in our framework by relating numerical quantities to the physical parameters describing time-dependent areal coverage in real-world filtration processes. First, we denote $F_{\textnormal{gr}}(t)$ (cells) to be the number of foulants that have grafted on the membrane at time $t$, and $F_{\textnormal{area}}$ ( $\text{\,}{\mathrm{m}}^{2}$ cell-1 ) to be the lengthwise cross-sectional area of a single foulant particle. Then $F_{\textnormal{gr}}(t)\cdot F_{\textnormal{area}}$ gives the areal coverage of the simulated membrane at time $t$ due to the accumulation of foulants.
As noted previously, we have $n_{exp}$ particles entering the channel at every time step, where $n_{exp}$ is a function of the foulant concentration and the initial flux. In practice, a typical $n_{exp}$ would be of the order of $10^{4}$. For computational efficiency in the simulation (where we would have to introduce $10^{4}$ at each time step), we scale $n_{exp}$ by $10^{\textnormal{s}}$ for some $s\in\mathbb{Z}^{+}$, and define the actual number of particles we introduce at each time step as $n=n_{exp}/10^{s}$ (cell  $\text{\,}{\mathrm{s}}^{-1}$). To compute the actual number of foulants that would attach to the membrane in the physical, real-world experiment, we scale $F_{\textnormal{gr}}(t)$ by $F_{\textnormal{conc}}/n$. Altogether, we model the areal membrane coverage due to the attached foulants at time $t$ per unit volume filtered ( $\text{\,}{\mathrm{m}}^{-1}\text{\,}\mathrm{s}$) as:
@@ 3.1.1 Model benchmarking
$F_{\textnormal{conc}}$ Foulant concentration (cells  $\text{\,}{\mathrm{m}}^{-3}$) $2.86\times 10^{12}$ $1.53\times 10^{13}$ $5.49\times 10^{12}$ $2.75\times 10^{12}$ [83], [19]
@@
Figure 4: Forward filtration simulations. Each simulation curve is the average of 100 runs. Parameters for each figure are given in Table 2. (a) Simulation of feed water with $2.86\times 10^{12}$ cells  $\text{\,}{\mathrm{m}}^{-3}$ of B. diminuta filtered through membrane with 400-nm pores at a constant TMP of $35.852\times 10^{3}\,$\text{\,}\mathrm{Pa}$\$ (5.2 psi) and initial flux of $3.48\times 10^{-5}\,$\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$\$. Experimental data from Figure 4 of [19]. (b) Simulation of feed water with $1.53\times 10^{13}$ cells  $\text{\,}{\mathrm{m}}^{-3}$ of B. diminuta filtered through membrane with 200-nm pores at a constant TMP of $33.577\times 10^{3}\,$\text{\,}\mathrm{Pa}$\$ (4.87 psi) and initial flux of $2.2\times 10^{-4}\,$\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$\$. Experimental data from Figure 3 of [83]. (c) Simulation of feed water with $5.49\times 10^{12}$ cells  $\text{\,}{\mathrm{m}}^{-3}$ S. marcescens filtered through membrane with 200-nm pores at a constant TMP of $27.165\times 10^{3}\,$\text{\,}\mathrm{Pa}$\$ (3.94 psi) and initial flux of $3.49\times 10^{-5}\,$\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$\$. Experimental data from Figure 7 of [19]. (d) Simulation of feed water with $2.75\times 10^{12}$ cells  $\text{\,}{\mathrm{m}}^{-3}$ of S. marcescens filtered through membrane with 400-nm pores at a constant TMP of $28.958\times 10^{3}\,$\text{\,}\mathrm{Pa}$\$ (4.2 psi) and initial flux of $1.72\times 10^{-4}\,$\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$\$. Experimental data from Figure 3 of [16].
@@
Consistent with the predictions of membrane fouling dominated by intermediate blocking [13, 83], the flux decline curves are concave up. As expected, experiments run with larger bacteria (Figure 4, bottom row) foul the membrane more quickly than those run with smaller bacteria (Figure 4, top row). Moreover, for the same bacteria type, a higher initial flux (Figure 4, right column) results in a steeper drop in flux over time. When comparing the results plotted in Figures 4a and 4c, we observe greater overall flux decline in the latter, where the bacteria size is larger and the membrane pore size is smaller, though the experiments start with similar initial fluxes. In Figures 4b and 4d, we see that for sufficiently high initial fluxes (an order of magnitude larger than Figures 4a and 4c) and large bacterial concentrations, the flux declines more much sharply. Across the experiments, a higher initial flux corresponds to higher Pe number values, which indicate that advection dominates over diffusion; this is reflected in the smoother flux decline curves of Figures 4b and 4d compared to Figures 4a and 4c.
@@
We rely primarily on five physical, real-world parameters to generate the flux decline simulations in Figure 4, as detailed in Table 2; the only model parameter fit to the data is the adjustment $A_{\textnormal{adj}}.$ Using these five physical parameters and fitting one parameter, our model performs very well by matching experimental flux decline values with residual errors of 0.431 (Figure 4a), 0.255 (Figure 4b), 0.830 (Figure 4c), and 0.676 (Figure 4d).
@@ 3.1.2 Model extensions
In Figure 5b, we run forward filtration simulations altering the initial flux. We compare the normalized flux decline shown in Figure 4b to simulations in which the initial flux is halved ($\cdot\cdot\cdot$) or doubled ($\cdot\cdot\cdot$). As the initial flux increases, so does the concavity of the flux decline curve, suggesting more rapid accumulation of foulants on the membrane.
@@ 3.2 Forward filtration-backwashing cycling
Table 3: In-code parameters and variables used for simulations in Figure 7. Most of the physical parameter values were derived or inferred from the original experiments [29, 28] or numerical experiments done by the same research group [21].
@@
n Input rate (cells  $\text{\,}{\mathrm{s}}^{-1}$) 118 -

### v2

@@ Abstract
Membrane filtration is widely used in water treatment to remove foulants from contaminated water. Foulant build-up on the membrane occludes the area open for fluid flow, which impairs the efficiency of the filtration operation by decreasing the flux through the membrane. Backwashing is a strategy to restore the membrane, wherein clean water is processed backward through the membrane to dislodge attached foulants. We develop a Monte Carlo model to simulate constant-pressure forward filtration and backwashing through dead-end, flat-sheet membranes, with membrane fouling dominated by intermediate blocking. We validate our model against real-world experiments conducted with different foulant types and concentrations and run under different filtration conditions. Relying primarily on measurable physical parameters and employing easy-to-implement parameter fitting techniques as needed, we show good agreement between experimental data and numerical simulations. We extend these results to predict flux behavior in forward filtration and backwashing when foulant properties or filtration conditions are varied. The newly developed model can be used to further investigate the impact of varying backwashing duration, frequency, and/or pressure on the rate of flux recovery.
@@ 1 Introduction
Increases in population, industry, and pollution over the past century have exacerbated the long-extant problem of clean water scarcity [13, 78, 80]. Water purification requires many stages of treatment, among which membrane filtration is favorable for its relatively low energy requirements and capital cost [27, 10]. This technology has been applied in many industries, including water treatment for wastewater [8, 75, 60], surface water [76, 37, 55], and groundwater [76, 10]; food and drink manufacturing [71, 16]; and downstream processing in the pharmaceutical industry [57, 77, 15, 63].
Membrane filtration operates by two main mechanisms: (1) physiochemical interactions between foulants and the membrane and (2) sieving [65, 11]. Our focus is on membrane microfiltration of fluid fouled by suspended solids, bacteria, and other undesired materials that are larger than the membrane pores [9]. Thus, here sieving is the dominant mechanism, and we split membrane filtration processes into two repeated steps: forward filtration (FF) followed by backwashing (BW). In forward filtration with dead-end, flat-sheet membranes, the fluid is pushed perpendicularly through the membrane and foulants accumulate directly onto its surface. The result is a phenomenon known as membrane fouling, wherein the membrane pores become increasingly blocked [72, 14]. When filtration is run under constant-pressure conditions, membrane fouling results in considerable flux decline, which, in the long run, impairs the efficiency of and decreases the yield from filtration procedures [33, 34]. To counteract this decline, operators employ “backwashing,” running clean water backward through the membrane to dislodge foulants that have attached to it. This process recovers some, though not all, of the flux lost during forward filtration.
@@ 2.2.1 Advection-Diffusion
Our Monte Carlo method builds on equation (3). We define the concentration of a given foulant in the contaminated water being filtered as $F_{\textnormal{conc}}$ (cells  $\text{\,}{\mathrm{m}}^{-3}$) and let $F_{0}$ be the number of foulants in the channel at time $t=0$ based on $F_{\textnormal{conc}}$ [50]. We compute $n_{\textnormal{exp}}$ (cell  $\text{\,}{\mathrm{s}}^{-1}$) as the number of foulants that enter the domain at every time step based on the initial flux, $J_{0}$ ( $\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$).
@@ 2.2.2 Foulant grafting & flux decline
When a foulant particle approaches the membrane, it either grafts to its surface or it bounces back into the channel flow. Whether each foulant particle attaches or bounces back is determined probabilistically. Any given foulant has an associated probability of grafting to the membrane, $P_{gr}$, that is uniformly distributed between 0 and some maximum probably of grafting $p_{gr}$, i.e., $P_{gr}\in\textnormal{Unif}(0,p_{gr})$. The probability $P_{gr}$ may be additionally adjusted by the particle size and whether other foulants are already stuck on the area of the membrane where a foulant would attach. We allow multiple foulants to occupy positions on the membrane in very close proximity to conceptualize foulant overlap.
We model fouling dynamics as dominated by “intermediate pore blocking” where foulants can overlap and cover both the membrane surface and the membrane pores due to their size [14, 84]; this creates a decline in flux over the course of forward filtration. Experimental analysis suggests that membrane filtration transitions from the intermediate pore blocking regime to the “cake filtration” regime when 10 – 20% of the membrane area remains open [84]. The rate of membrane area coverage depends on the contamination of the water, the membrane pore sizes, and the duration of filtration. At the concentration levels, pore sizes, and filtration runtimes modeled in the experimental set-ups we compare to in Section 3.1, this degree of membrane area coverage is not achieved and, thus, the cake formation phase of filtration is not reached [84].
@@
We adapt equation (4) to model the flux decline in our framework by relating numerical quantities to the physical parameters describing time-dependent areal coverage in real-world filtration processes. First, we denote $F_{\textnormal{gr}}(t)$ (cells) to be the number of foulants that have grafted on the membrane at time $t$, and $F_{\textnormal{area}}$ ( $\text{\,}{\mathrm{m}}^{2}$ cell-1 ) to be the lengthwise cross-sectional area of a single foulant particle. Then $F_{\textnormal{gr}}(t)\cdot F_{\textnormal{area}}$ gives the areal coverage of the simulated membrane at time $t$ due to the accumulation of foulants. We note that such areal coverage is likely an overestimation due to partial overlap of some grafted foulants during intermediate blocking, which is permitted in the model, as described earlier.
As noted previously, we have $n_{\textnormal{exp}}$ particles entering the channel at every time step, where $n_{\textnormal{exp}}$ is a function of the foulant concentration and the initial flux. In practice, a typical $n_{\textnormal{exp}}$ would be of the order of $10^{4}$. For computational efficiency in the simulation (where we would have to introduce $10^{4}$ at each time step), we scale $n_{\textnormal{exp}}$ by $10^{\textnormal{s}}$ for some $s\in\mathbb{Z}^{+}$, and define the actual number of particles we introduce at each time step as $n=n_{\textnormal{exp}}/10^{s}$ (cell  $\text{\,}{\mathrm{s}}^{-1}$). To compute the actual number of foulants that would attach to the membrane in the physical, real-world experiment, we scale $F_{\textnormal{gr}}(t)$ by $F_{\textnormal{conc}}/n$. Altogether, we model the areal membrane coverage due to the attached foulants at time $t$ per unit volume filtered ( $\text{\,}{\mathrm{m}}^{-1}\text{\,}\mathrm{s}$) as:
@@ 3.1.1 Model benchmarking
Consistent with the predictions of membrane fouling dominated by intermediate blocking [14, 84], the flux decline curves are concave up on long time-scales. The curves in Figures 4a and 4c are initially concave down due to the lower initial flux, $J_{0}$, employed in these experiments. Lower fluxes translate to slower arrival of foulants to the membrane and, thus, slower onset of the upward concavity characteristic of intermediate blocking.
@@
$F_{\textnormal{conc}}$ Foulant concentration (cells  $\text{\,}{\mathrm{m}}^{-3}$) $2.86\times 10^{12}$ $1.53\times 10^{13}$ $2.75\times 10^{12}$ $5.49\times 10^{12}$ [84], [20]
@@
Figure 4: Forward filtration simulations. Each simulation curve is the average of 100 runs. Parameters for each figure are given in Table 2. (a) Simulation of feed water with $2.86\times 10^{12}$ cells  $\text{\,}{\mathrm{m}}^{-3}$ of B. diminuta filtered through membrane with 400-nm pores at a constant TMP of $35.852\times 10^{3}\,$\text{\,}\mathrm{Pa}$\$ (5.2 psi) and initial flux of $3.48\times 10^{-5}\,$\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$\$. Experimental data from Figure 4 of [20]. (b) Simulation of feed water with $1.53\times 10^{13}$ cells  $\text{\,}{\mathrm{m}}^{-3}$ of B. diminuta filtered through membrane with 200-nm pores at a constant TMP of $33.577\times 10^{3}\,$\text{\,}\mathrm{Pa}$\$ (4.87 psi) and initial flux of $2.2\times 10^{-4}\,$\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$\$. Experimental data from Figure 3 of [84]. (c) Simulation of feed water with $2.75\times 10^{12}$ cells  $\text{\,}{\mathrm{m}}^{-3}$ S. marcescens filtered through membrane with 200-nm pores at a constant TMP of $27.165\times 10^{3}\,$\text{\,}\mathrm{Pa}$\$ (3.94 psi) and initial flux of $3.49\times 10^{-5}\,$\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$\$. Experimental data from Figure 7 of [20]. (d) Simulation of feed water with $5.49\times 10^{12}$ cells  $\text{\,}{\mathrm{m}}^{-3}$ of S. marcescens filtered through membrane with 400-nm pores at a constant TMP of $28.958\times 10^{3}\,$\text{\,}\mathrm{Pa}$\$ (4.2 psi) and initial flux of $1.72\times 10^{-4}\,$\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$\$. Experimental data from Figure 3 of [17].
@@
As expected, experiments run with larger bacteria (Figure 4, bottom row) foul the membrane more quickly than those run with smaller bacteria (Figure 4, top row). Moreover, for the same bacteria type, a higher initial flux (Figure 4, right column) results in a steeper drop in flux over time. When comparing the results plotted in Figures 4a and 4c, we observe greater overall flux decline in the latter, where the bacteria size is larger and the membrane pore size is smaller, though the experiments start with similar initial fluxes. In Figures 4b and 4d, we see that for sufficiently high initial fluxes (an order of magnitude larger than Figures 4a and 4c) and large bacterial concentrations, the flux declines more much sharply. Across the experiments, a higher initial flux corresponds to higher Pe number values, which indicate that advection dominates over diffusion; this is reflected in the smoother flux decline curves of Figures 4b and 4d compared to Figures 4a and 4c.
@@
We rely primarily on five physical, real-world parameters to generate the flux decline simulations in Figure 4, as detailed in Table 2; the only model parameter fit to the data is the adjustment $A_{\textnormal{adj}}.$ Using these five physical parameters and fitting one parameter, our model performs very well by matching experimental flux decline values with residual errors of 0.432 (Figure 4a), 0.256 (Figure 4b), 0.835 (Figure 4c), and 0.676 (Figure 4d).
@@ 3.1.2 Model extensions
In Figure 5b, we run forward filtration simulations altering the initial flux. We compare the normalized flux decline shown in Figure 4b to simulations in which the initial flux is halved ($\cdot\cdot\cdot$) or doubled ($\cdot\cdot\cdot$). As the initial flux increases, so does the concavity of the flux decline curve, suggesting more rapid accumulation of foulants on the membrane. Conversely, when halving the flux, initial concave down behavior appears due to the longer time needed for foulants to reach and begin accumulating on the membrane.
@@ 3.2 Forward filtration-backwashing cycling
Table 3: In-model parameters and variables used for simulations in Figure 7. Most of the physical parameter values were derived or inferred from the original experiments [30, 29] or numerical experiments done by the same research group [22].
@@
n Input rate (cells  $\text{\,}{\mathrm{s}}^{-1}$) 15 -